OpenStudy (anonymous):

Can some one please give me directions on how this g(x) = ex -4 Description of transformation: Equation(s) for the Horizontal Asymptote(s): y-intercept in (x, y) form

OpenStudy (anonymous):

The equation of the horizontal asymptote for the equation\[f(x)=e^x\]is the line y=0. So if we shift the graph down by 4 units, we get the equation \[g(x)=e^x-4\]Every point on the entire graph is shifted down by 4 units, which means our horizontal asymptote is now at y=-4. To find the y-intercept for any equation, let x=0 and solve for y. \[y=e^0-4\]\[y=1-4=3\]This gives the coordinates (0,-3) for the y-intercept.